OpenStudy (anonymous):

what is the value of x in this equation: 5-3(x+1)=6?

OpenStudy (anonymous):

5-3x-3=6 so 5-3-6=3x ---> -4=3x x=-4/3

OpenStudy (anonymous):

that is not any of the answers available. There is 0. 8/3, -1,-2

OpenStudy (anonymous):

are you sure you wrote the question correctly? because when i substitute -4/3 back into the equation, i get 6

OpenStudy (anonymous):

thanks it is 6-x so x is what?

OpenStudy (anonymous):

you mean as in the equation is equal to 6-x?

OpenStudy (anonymous):

in that case, x would be -2

OpenStudy (anonymous):

5-3(x+1)=6-x then x is: 0, -1,-2, 8/3 this is the right way

OpenStudy (anonymous):

thanks for helping i appreciate it have a great weekend.